OpenStudy (anonymous):

write the equation for the tangent line in slope intercept form 2x^2+3xy+3y^2=17;(2,1)

OpenStudy (anonymous):

anything will help please!

OpenStudy (anonymous):

did you get \[y'\]?

OpenStudy (anonymous):

i got it wrong lol thats what happened..ive done it twice..i only get one more chance

OpenStudy (anonymous):

ok lets see \[2x^2+3xy+3y^2=17\] \[4x+3y+3xy'+6yy'=0\] right?

OpenStudy (anonymous):

i got that yes

OpenStudy (anonymous):

\[4\times 2+3\times 1+3\times 2y'+6\times 1\times y'=0\] \[8+3+6y'+6y'=0\] \[11+12y'=0\] \[12y'=-11\] \[y=-\frac{11}{12}\] check my algebra since you only have one more chance

OpenStudy (anonymous):

well i need the actual equation..so it would be y=-11/12x with the y intercept

OpenStudy (anonymous):

assuming i did not make an arithmetic mistake, which you really should check, you have the slope as \[-\frac{11}{12}\]and then point - slope formula give \[y-1=-\frac{11}{12}(x-2)\]

OpenStudy (anonymous):

ill check it...but is it going to be -11/12x-1?

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

it will be what i wrote, and if you solve for y you should get \[y=-\frac{11}{12}x+\frac{17}{6}\]

OpenStudy (anonymous):

ok ill check the math and let you know

OpenStudy (anonymous):

it was right!!! =)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw don't forget point slope formula!